If your measurement of some quantity is uncertain
Answer #1


area = L^2 = (10.5*10.5)


area = 110.25


percent uncertainity = (5% + 5%)

obsolute uncertainity = 0.1*110.25 = 11.025 in^2


area = 110.25 +/- 11.025


maximum area = 110.25 + 11.025 = 121 in^2

minimum area = 110.25-11.025 = 99 in^2

area = 110.25 +/- 11.025